1. Calculate the freezing and boiling points of a solution prepared by dissolving 15.5g of AI(NO3)3 in 200.0g of water. (Molar mass of AI(NO3)3 is 212.996 g/mol).


2. A solution is prepared by dissolving 120 grams of NaCI in 450 grams of water. Find the freezing and boiling point of this solution. (Molar mass of NaCI is 58.44 g/mol).


1
Expert's answer
2021-04-30T04:51:05-0400

1. Molality = Mole/kg = Mass/(molar mass × mass in kg)

= 15.5/(212.996 ×0.2) = 0.36

Tsoln = TWater + k.m.i

= 100 + 0.51(0.36)(4)

= 100.73°C (boiling point)

Freezing point

Tsoln = Tice + k.m.i

= 0 +(-1.86)(0.36)(4)

= -2.68°C


2. Molality = Mole/kg = Mass of solute /(molar mass × mass of solvent in kg)

120/(58.44×0.45) = 4.56

Tsoln = TWater + k.m.i

= 100°C + 0.51(4.56)(2)

= 104.65°C (boiling point)


Freezing point

Tsoln = Tice + k.m.i

= 0° + (-1.86)(4.56)(2)

= -16.96°C